# ARR: Question Answering with Large Language Models via Analyzing, Retrieving, and Reasoning

### v2

@@ Abstract
Large language models (LLMs) achieve remarkable performance on challenging benchmarks that are often structured as multiple-choice question-answering (QA) tasks. Zero-shot Chain-of-Thought (CoT) prompting enhances reasoning in LLMs but provides only vague and generic guidance (“think step by step”). This paper introduces ARR, an intuitive and effective zero-shot prompting method that explicitly incorporates three key steps in QA solving: analyzing the intent of the question, retrieving relevant information, and reasoning step by step. Comprehensive experiments across diverse and challenging QA tasks demonstrate that ARR consistently improves the Baseline (without ARR prompting) and outperforms CoT. Ablation and case studies further validate the positive contributions of each component: analyzing, retrieving, and reasoning. Notably, intent analysis plays a vital role in ARR. Additionally, extensive evaluations across various model sizes, LLM series, and generation settings solidify the effectiveness, robustness, and generalizability of ARR. [^1] Source code: https://github.com/YuweiYin/ARR
@@ 1 Introduction
Recent advancements have introduced various methods to enhance LLM reasoning abilities Qiao et al. (2023); Sun et al. (2023), with Chain-of-Thought (CoT) prompting proving effective across various tasks Li et al. (2024). Key variants include few-shot CoT Wei et al. (2022), which provides rationale-based exemplars for in-context learning Brown et al. (2020); Dong et al. (2024), and zero-shot CoT Kojima et al. (2022), which employs general instructions such as “Let’s think step by step.” Due to its simplicity and effectiveness, zero-shot CoT has gained extensive adoption. By appending this trigger sentence to the original QA prompt, LLMs generate step-by-step reasoning to improve question-answering performance.
@@
Despite its widespread use, zero-shot CoT prompting provides only generic reasoning guidance. As illustrated in Figure 1, answering complex questions typically involves three key steps: (1) analyzing the question’s intent Adams (1986); Mele (1989); Mele and Moser (1994) to obtain a thorough context understanding, a clear problem-solving target, and a purposeful planning guide, (2) retrieving relevant information from context, external sources, or memory for supportive reference Jones and Steinhardt (2022); Shi et al. (2023), and (3) systematically applying inductive and deductive reasoning Clark (1969); Johnson-Laird (1999); Heit (2000); Hayes and Heit (2018). Therefore, we hypothesize that an effective prompt should direct LLMs to complete these steps. To verify this hypothesis, we propose a refined zero-shot prompting method, ARR, which explicitly incorporates these three elements: Analyzing, Retrieving, and Reasoning. Specifically, ARR employs the answer trigger sentence: “Let’s analyze the intent of the question, find relevant information, and answer the question with step-by-step reasoning.” This structured approach is expected to enhance the performance across diverse QA tasks and various models, akin to the improvements observed with zero-shot CoT upon its first introduction Kojima et al. (2022).
@@
To evaluate the effectiveness of the proposed ARR method, we test the performance (accuracy) of open-weights LLMs Dubey et al. (2024) on 10 multiple-choice QA datasets, covering reading comprehension Clark et al. (2019); Liu et al. (2020), commonsense reasoning Talmor et al. (2019); Sap et al. (2019), world knowledge Welbl et al. (2017); Mihaylov et al. (2018); Clark et al. (2018), and multitask understanding Suzgun et al. (2023); Hendrycks et al. (2021); Wang et al. (2024b). Compared to the Baseline method without a specific trigger sentence and the zero-shot CoT method with general prompts, ARR consistently improves QA performance across all datasets, demonstrating its effectiveness and superiority. Additionally, ablation studies show that each individual component of ARR–Analyzing, Retrieving, and Reasoning–outperforms both the Baseline and CoT methods, confirming their positive contributions. Notably, the Analyzing-only setting yields the largest performance gain on average, highlighting the critical role of intent analysis in question answering. Beyond quantitative results, we provide qualitative case studies to reveal problems in the Baseline and CoT methods such as intent misunderstanding, context misuse, and faulty reasoning.
@@
Furthermore, we conduct extensive experiments across various settings to assess the generalizability of the proposed ARR method. ARR consistently outperforms alternatives across different model sizes, LLM series (architectures), generation temperatures, and few-shot scenarios. These comprehensive experiments and analyses further solidify its effectiveness, robustness, and adaptability. The key contributions of this work are as follows:
@@
    This paper proposes ARR, an intuitive, general, and effective zero-shot prompting method to improve LLM performance in various question-answering tasks.
@@
    Comprehensive experiments across diverse QA tasks demonstrate that ARR consistently outperforms the Baseline and CoT methods. Ablation and case studies further validate the positive contributions of each component.
@@ 2.1 LLM Prompting
Recent large language models (LLMs) Dubey et al. (2024); Lambert et al. (2024); Liu et al. (2024) are pre-trained on large-scale text corpora curated from the Internet Soldaini et al. (2024); Penedo et al. (2024); Weber et al. (2024). Their advanced text understanding and generation capabilities Hurst et al. (2024); Anthropic (2024); Team et al. (2024a) have significantly revolutionized the field of natural language processing (NLP). Consequently, the NLP paradigm is shifting toward a framework comprising pre-training, post-training, and prompting Liu et al. (2023), with post-training focusing on aligning models with human preferences Ouyang et al. (2022); Bai et al. (2022); Rafailov et al. (2023) rather than fine-tuning for specific downstream tasks Devlin et al. (2019). After the training stages, LLMs can generate satisfactory responses to natural language instructions and questions, highlighting the growing importance of prompt design White et al. (2023); Giray (2023); Sahoo et al. (2024). In this work, we propose an intuitive, general, and effective prompting method to enhance LLM performance in question-answering.
@@ 2.2 LLM Reasoning
Recent LLM research increasingly emphasizes reasoning abilities Qiao et al. (2023); Sun et al. (2023). Chain-of-Thought (CoT) is a prompting strategy that enhances problem-solving by guiding LLMs to generate intermediate reasoning steps. Main variants include zero-shot CoT Kojima et al. (2022) that uses general instructions such as “Let’s think step by step” and few-shot CoT Wei et al. (2022) that provides exemplars with rationales to leverage in-context learning Brown et al. (2020); Dong et al. (2024). Building on CoT, various reasoning techniques have emerged Zhou et al. (2023a, b); Wang et al. (2023a); Yasunaga et al. (2024); Wang and Zhou (2024). Some studies explore optimal reasoning paths through self-consistency Wang et al. (2023c); Chen et al. (2023) or tree-like searches Yao et al. (2023), while others investigate self-refinement Madaan et al. (2023), self-correction Huang et al. (2024); Tyen et al. (2024); Chen et al. (2024), self-verification Cobbe et al. (2021); Li et al. (2023b); Lightman et al. (2024), and self-evolution Guan et al. (2025); Lee et al. (2025) mechanisms. Beyond prompting and generation-based approaches, post-training methods Chu et al. (2025), particularly those leveraging reinforcement learning (RL) Sutton and Barto (2018), have been developed to enhance reasoning capabilities Shao et al. (2024); Wang et al. (2024a); Setlur et al. (2024); Xu et al. (2025). As a reasoning-eliciting prompting approach, ARR effectively complements existing research by guiding LLMs through three essential steps: intent analysis, information retrieval, and step-by-step reasoning.
@@ 3.1 Question Answering Data
In this work, we consider multiple-choice question-answering (MCQA) tasks with one correct answer, where the model is asked to answer the question by selecting an option from a list of choices. Formally, let $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ be a MCQA dataset, where $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n}\}$ is the input information, $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$ is the corresponding correct-choice label ($y_{i}\in\mathbb{R}$), and $n$ is the number of instances in $\mathcal{D}$.
@@
The answer trigger sentence $\phi$ is the only difference between the proposed ARR method and baseline methods in each experiment. Figure 2 presents each $\phi$ used in Baseline, zero-shot CoT, and our ARR methods. For simplicity, $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$ is used in the rest of the paper for both open- and closed-book QA.
@@ 3.2 Multiple-Choice Question Answering
Stage 1: Reasoning Generation.
Let $\tilde{x}_{i}$ be the tokenized representation of text $x_{i}$. The decoder-only Transformer-based Vaswani et al. (2017); Radford et al. (2018) LLM $\mathcal{M}$ takes $\tilde{x}_{i}$ as input and generate a new token after each timestep. The model freely generates the text response given by
@@
Stage 2: Option Selection.
@@ 4.1 Datasets
Table 2: Main experiments. The zero-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on various multiple-choice QA datasets using different answer trigger sentences $\phi$. (1) w/o Reason: directly selecting an option without rationales; (2) Baseline: $\phi$ is “Answer:”; (3) CoT Kojima et al. (2022): $\phi$ is “Answer: Let’s think step by step.”; (4) ARR: our method that elicits intent analysis, information retrieval, and step-by-step reasoning.
@@ 4.1.4 Multitask Understanding
BIG-Bench Hard Suzgun et al. (2023) is a suite challenging tasks filtered from BIG-Bench Srivastava et al. (2023). Solving these problems often requires multi-step reasoning. In this work, $4$ (out of $27$) subtasks in BBH are discarded as they are not multiple-choice QA tasks.
@@ 4.2 Models
Our experiments adopt open-weights, decoder-only, and Transformer-based Vaswani et al. (2017) LLMs. We mainly employ LLaMA3-8B-Chat Dubey et al. (2024), an instruction-following LLM with $8$ billion model parameters, and use the model implementation and checkpoints provided by Hugging Face Transformers Wolf et al. (2020). In generalizability experiments, we also explore LLaMA3-Chat models of different sizes in § 6.1 and 7B-Chat models of different LLM series in § 6.2, i.e., Qwen2.5 Yang et al. (2024), Gemma Team et al. (2024b, c), and Mistral Jiang et al. (2023).
@@ A R R BoolQ LogiQA CSQA SIQA SciQ OBQA ARC BBH MMLU MMLU-Pro
Table 4: Ablation study results. The accuracy scores (%) of the LLaMA3-8B-Chat model on various multiple-choice QA datasets using different answer trigger sentences $\phi$ (Analyzing, Retrieving, and Reasoning).
@@ 5.1 QA Performance
The main experiments test the zero-shot QA performance of LLaMA3-8B-Chat Dubey et al. (2024) on various multiple-choice QA datasets. The only difference between Baseline, zero-shot CoT Kojima et al. (2022), and ARR is the answer trigger sentence $\phi$ shown in Figure 2. The results in Table 2 demonstrate that our ARR method boosts the Baseline method by a large margin, with an improvement of $+4.1\%$ on average. In addition, ARR consistently outperforms zero-shot CoT prompting across all QA datasets, highlighting its universal superiority in various task types including reading comprehension, commonsense reasoning, world knowledge, and multitask understanding. Moreover, the “w/o Reason” method, which directly selects options without relying on rationales ($r_{i}$ in Eq. 2), performs significantly worse, emphasizing the benefits of our two-stage QA approach.
@@ 5.2 Ablation Study
To better understand the performance gains shown in Table 2, we conduct an ablation study to explore the efficacy of each component of the ARR method, i.e., analyzing, retrieving, and reasoning. Specifically, we test the model’s QA performance using the five different answer trigger sentences $\phi$ in Table 3. Table 4 reports the accuracy scores of LLaMA3-8B-Chat under different ablation cases, where ➀ is the full version of ARR and ➄ is equivalent to the “Baseline” method in Table 2. In ➁, ➂, and ➃, $\phi$ only contains one single component, i.e., analyzing, retrieving, and reasoning, respectively.
@@
We observe that all the single-component ARR settings (➁, ➂, and ➃) outperform the Baseline method (➄) by a large margin, which verifies that each ARR component contributes positively. In addition, ➃ outperforms the CoT method in Table 2, which means the “Reasoning” prompting of our ARR method (i.e., “answer the question with step-by-step reasoning”) is better than zero-shot CoT prompting (i.e., “think step by step”) for QA tasks. Furthermore, the complete ARR method (➀) has a higher accuracy score than the Retrieving-only (➂) and Reasoning-only (➃) methods, meaning the intent analysis benefits the other two “R” parts.
Notably, the Analyzing-only setting (➁) brings the greatest improvement gain, suggesting the significance of analyzing the intent of the question. ➁ even outperforms the full ARR version (➀) on average, mainly because it performs excellently on the OBQA dataset (86.80%), which means understanding the question’s intent in this task is vital. Excluding OBQA, the full ARR method achieves an average accuracy of 68.41%, surpassing ➁ (68.27%). Hence, we dive into a concrete case study to examine the impact of intent analysis on QA solving.
Task Input $x_{i}$ Fact: the summer solstice is on June 21st in the northern hemisphere. Question: The summer solstice in the northern hemisphere is four months before (A) May   (B) July   (C) April   (D) October
Baseline $\phi_{i}$ Answer:
Reasoning $r_{i}$ (B) July. Explanation: The summer solstice is the longest day of the year and marks the beginning of summer. It is four months before the beginning of winter. The beginning of winter is in December, so the summer solstice is in June. Therefore, the summer solstice is four months before July.
CoT $\phi_{i}$ Answer: Let’s think step by step.
Reasoning $r_{i}$ The summer solstice is on June 21st. If we count four months before that, we get March 21st. March is the third month of the year, so the answer is (C). April is the fourth month, May is the fifth month, and October is the tenth month. None of these months is four months before June. July is the seventh month, which is five months after June, so it is not the correct answer either. The correct answer is (C).
Reasoning $r_{i}$ The question asks for the month that is four months after the summer solstice. The summer solstice is on June 21st. Four months after June 21st is October 21st. Therefore, the correct answer is (D) October.
Table 5: Case study. An instance from the OpenBookQA dataset, where our ARR method correctly answers the question but the Baseline and CoT methods fail because of intent misunderstanding or faulty reasoning.
5.3 Case Study
Table 5 shows a case in the OBQA dataset Mihaylov et al. (2018) drawn from the main experiments [^3] Please refer to Appendix B for more cases., where our ARR method correctly answers the question but the Baseline and CoT methods fail. We can observe that the baseline method’s reasoning is incorrect: After stating that “the summer solstice is in June,” it wrongly concludes that “the summer solstice is four months before July.” The CoT method misunderstands the question, resulting in counting four months before June instead of after June. In contrast, our ARR method identifies the question’s intent clearly, leading to a correct reasoning path and final answer.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities on complex evaluation benchmarks, many of which are formulated as question-answering (QA) tasks. Enhancing the performance of LLMs in QA contexts is becoming increasingly vital for advancing their development and applicability. This paper introduces ARR, an intuitive, effective, and general QA solving method that explicitly incorporates three key steps: analyzing the intent of the question, retrieving relevant information, and reasoning step by step. Notably, this paper is the first to introduce intent analysis in QA, which plays a vital role in ARR. Comprehensive evaluations across 10 diverse QA tasks demonstrate that ARR consistently outperforms the baseline methods. Ablation and case studies further validate the positive contributions of each ARR component. Furthermore, experiments involving variations in prompt design indicate that ARR maintains its effectiveness regardless of the specific prompt formulation. Additionally, extensive evaluations across various model sizes, LLM series, and generation settings solidify the effectiveness, robustness, and generalizability of ARR. [^1] Source code: https://github.com/YuweiYin/ARR
@@ 1 Introduction
As illustrated in Figure 1, answering complex questions typically involves three key steps: (1) analyzing the question’s intent Adams (1986); Mele (1989); Mele and Moser (1994) to obtain a thorough context understanding, a clear problem-solving target, and a purposeful planning guide, (2) retrieving relevant information from context, external sources, or memory for supportive reference Jones and Steinhardt (2022); Shi et al. (2023), and (3) systematically applying inductive and deductive reasoning Clark (1969); Johnson-Laird (1999); Heit (2000); Hayes and Heit (2018). Therefore, we hypothesize that an effective solution should direct LLMs to complete these key steps. To verify this hypothesis, we propose a refined QA framework, ARR, which explicitly incorporates these three elements: Analyzing, Retrieving, and Reasoning.
As a general framework, ARR can be implemented by simply prompting LLMs to follow the three steps or more elaborately by collecting such three-step data for LLM training. In this work, we investigate ARR as a test-time prompting method because this is the most natural and direct approach to assess its effectiveness when applied to pre-trained foundation models like LLMs. Specifically, ARR adopts the following answer trigger sentence at the beginning of LLM’s output: “Let’s analyze the intent of the question, find relevant information, and answer the question with step-by-step reasoning.” This explicit and structured approach is expected to enhance the performance across diverse QA tasks and various models.
@@
To evaluate our ARR method, we test the performance (accuracy) of open-weights LLMs Dubey et al. (2024) on 10 diverse QA datasets, covering reading comprehension Clark et al. (2019); Liu et al. (2020), commonsense reasoning Talmor et al. (2019); Sap et al. (2019), world knowledge Welbl et al. (2017); Mihaylov et al. (2018); Clark et al. (2018), and multitask understanding Suzgun et al. (2023); Hendrycks et al. (2021); Wang et al. (2024b). Compared to the Direct Answer (DA) method without a specific trigger sentence and zero-shot Chain-of-Thought (CoT) method Kojima et al. (2022) with a generic prompt (“Let’s think step by step.”), ARR consistently improves QA performance across all datasets, demonstrating its effectiveness and superiority. Additionally, ablation studies show that each component of ARR (Analyzing, Retrieving, and Reasoning) outperforms the baselines, confirming their individual positive contributions. Notably, Intent Analysis—first introduced by ARR—yields the largest performance gain on average, highlighting the critical role of intent analysis in question answering. Moreover, experiments on five distinct prompt variants—each representing a paraphrased version of the original ARR prompt—demonstrate that ARR consistently remains effective irrespective of the particular prompt design.
@@
Furthermore, we conduct extensive experiments across various settings to assess the generalizability of our method. ARR consistently outperforms alternatives across different model sizes, LLM series (architectures), generation temperatures, and few-shot scenarios. These comprehensive experiments and analyses further solidify its effectiveness, robustness, and adaptability. Beyond quantitative results, we provide case studies (Appendix C) to reveal problems in the baselines such as intent misunderstanding, context misuse, and faulty reasoning. The key contributions of this work are as follows:
@@
    We propose ARR, an intuitive, effective, and general QA framework of three key components: intent analysis, information retrieval, and logical reasoning.
@@
    Comprehensive experiments across diverse QA tasks demonstrate that ARR consistently outperforms baseline methods. Ablation and case studies further validate the positive contributions of each component.
@@ 2.1 LLM Prompting
Recent large language models (LLMs) Dubey et al. (2024); Lambert et al. (2024); Liu et al. (2024) are pre-trained on large-scale text corpora curated from the Internet Soldaini et al. (2024); Penedo et al. (2024); Weber et al. (2024). Their advanced text understanding and generation capabilities Hurst et al. (2024); Anthropic (2024); Team et al. (2024a) have significantly revolutionized the field of natural language processing (NLP). Consequently, the NLP paradigm is shifting toward a framework comprising pre-training, post-training, and prompting Liu et al. (2023), with post-training focusing on aligning models with human preferences Ouyang et al. (2022); Bai et al. (2022); Rafailov et al. (2023) rather than fine-tuning for specific downstream tasks Devlin et al. (2019). After the training stages, LLMs can generate satisfactory responses to natural language instructions and questions, highlighting the growing importance of prompt design White et al. (2023); Giray (2023); Sahoo et al. (2024). In this work, we implement ARR in a prompting manner to empower LLMs.
@@ 2.2 LLM Reasoning
Recent LLM research increasingly emphasizes reasoning abilities Qiao et al. (2023); Sun et al. (2023). Chain-of-Thought (CoT) Kojima et al. (2022); Wei et al. (2022) is a prompting strategy that enhances problem-solving by guiding LLMs to generate intermediate reasoning steps. Building on CoT, various reasoning techniques have emerged Zhou et al. (2023a, b); Wang et al. (2023a); Yasunaga et al. (2024); Wang and Zhou (2024). Some studies explore optimal reasoning paths through self-consistency Wang et al. (2023c); Chen et al. (2023) or tree-like searches Yao et al. (2023), while others investigate self-refinement Madaan et al. (2023), self-correction Huang et al. (2024); Tyen et al. (2024); Chen et al. (2024), self-verification Cobbe et al. (2021); Li et al. (2023b); Lightman et al. (2024), and self-evolution Guan et al. (2025); Lee et al. (2025) mechanisms. Beyond prompting and generation-based approaches, post-training methods Chu et al. (2025), particularly those leveraging reinforcement learning (RL) Sutton and Barto (2018), have been developed to enhance reasoning capabilities Shao et al. (2024); Wang et al. (2024a); Setlur et al. (2024); Xu et al. (2025). As a structured reasoning-enhancing method, ARR effectively complements existing research by guiding LLMs through three essential steps: intent analysis, information retrieval, and logical reasoning.
@@ 3.1 Question Answering Data
In this work, we consider multiple-choice question-answering (MCQA) tasks with one correct answer, where the model is asked to answer the question by selecting an option from a list of choices. Formally, let $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ be an MCQA dataset, where $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n}\}$ is the input information, $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$ is the corresponding correct-choice label ($y_{i}\in\mathbb{R}$), and $n$ is the number of instances in $\mathcal{D}$.
@@
The answer trigger sentence $\phi$ is the only difference between the proposed ARR method and baseline methods in each experiment. Figure 2 presents each $\phi$ used in Direct Answer (DA), zero-shot CoT, and our ARR methods. For simplicity, $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$ is used in the rest of the paper for both open- and closed-book QA.
@@ 3.2 Multiple-Choice Question Answering
Stage 1: Reasoning Generation (RG).
Let $\tilde{x}_{i}$ be the tokenized representation of text $x_{i}$. The decoder-only Transformer-based Vaswani et al. (2017); Radford et al. (2018) LLM $\mathcal{M}$ takes $\tilde{x}_{i}$ as input and generate a new token after each timestep. The model freely generates the text response by
@@
Stage 2: Option Selection (OS).
@@ 4.1 Datasets
Table 2: Main experiments. The zero-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on various QA benchmarks using different answer trigger sentences $\phi$. (1) w/o RG: directly selecting an option without Reasoning Generation; (2) DA (Direct Answer): ϕ=italic-$\phi=$Answer:”; (3) CoT Kojima et al. (2022): ϕ=italic-$\phi=$Answer: Let’s think step by step.”; (4) ARR: our method that elicits intent analysis, information retrieval, and logical reasoning.
@@ 4.1.4 Multitask Understanding
BIG-Bench Hard Suzgun et al. (2023) is a suite challenging tasks filtered from BIG-Bench Srivastava et al. (2023). Solving these problems often requires multi-step reasoning.
@@ 4.2 Models
Our experiments adopt open-weights, decoder-only, and Transformer-based Vaswani et al. (2017) LLMs. We mainly employ LLaMA3-8B-Chat Dubey et al. (2024), an instruction-following LLM with $8$ billion model parameters, and use the model implementation provided by Hugging Face Transformers Wolf et al. (2020). In generalizability experiments, we also explore LLaMA3-Chat models of different sizes in § 6.1 and 7B-Chat models of different LLM series in § 6.2, i.e., Qwen2.5 Yang et al. (2024), Gemma Team et al. (2024b, c), and Mistral Jiang et al. (2023).
@@ A R R BoolQ LogiQA CSQA SIQA SciQ OBQA ARC BBH MMLU MMLU-Pro
Table 4: Ablation study results. The accuracy scores (%) of the LLaMA3-8B-Chat model on diverse QA datasets using different answer trigger sentences $\phi$ (Analyzing, Retrieving, and Reasoning).
@@ 5.1 QA Performance
The main experiments test the zero-shot QA performance of LLaMA3-8B-Chat Dubey et al. (2024) on diverse QA datasets. The only difference between Direct Answer (DA), zero-shot CoT Kojima et al. (2022), and ARR is the answer trigger sentence $\phi$ shown in Figure 2. The results in Table 2 demonstrate that our ARR method boosts the DA method by a large margin, with an improvement of $+4.1$ points on average. In addition, ARR consistently outperforms zero-shot CoT prompting across all QA datasets, highlighting its universal superiority in various task types including reading comprehension, commonsense reasoning, world knowledge, and multitask understanding [^3] Please refer to Appendix C for detailed case studies.. Moreover, the “w/o RG” method, which directly selects options without relying on rationales ($r_{i}$ in Eq. 2), performs significantly worse, emphasizing the benefits of our two-stage QA approach.
@@ 5.2 Ablation Study
To better understand the performance gains shown in Table 2, we conduct an ablation study to explore the efficacy of each component of the ARR method, i.e., analyzing, retrieving, and reasoning. Specifically, we test the model’s QA performance using the five different answer trigger sentences $\phi$ in Table 3. Table 4 reports the accuracy scores of LLaMA3-8B-Chat under different ablation cases, where ➀ is the full version of ARR and ➄ is equivalent to the “DA” method in Table 2. In ➁, ➂, and ➃, $\phi$ only contains one single component, i.e., analyzing, retrieving, and reasoning, respectively.
@@
We observe that all the single-component ARR settings (➁, ➂, and ➃) outperform the DA method (➄) by a large margin, which verifies that each ARR component contributes positively. Furthermore, the complete ARR method (➀) has a higher accuracy score than the Retrieving-only (➂) and Reasoning-only (➃) methods, meaning the intent analysis benefits the other two “R” parts. Notably, the Intent Analysis component (➁) brings the greatest improvement gain, suggesting the significance of analyzing the intent of the question.
Observing Table 4, each of ARR-full (➀), Analyzing-only (➁), Retrieving-only (➂), and Reasoning-only (➃) settings outperforms CoT and and DA in Table 2. Moreover, Analyzing-only (70.12%) and Retrieving-only (69.12%) settings (i.e., excluding the “Reasoning” component) beat CoT (68.14%), showing the fundamental difference between our ARR method and existing prompting methods. The experimental results shed light on future exploration on incorporating ARR steps, especially intent analysis, into problem solving.
5.3 Prompt Variants
Figure 3: Experiments on prompt variants. The average performance (Accuracy %) of the LLaMA3-8B-Chat model on 10 QA datasets using different ARR prompt variants (“V1”–“V5”).
To demonstrate that ARR works effectively irrespective of specific prompt design, we conduct experiments on different ARR prompt variants. The original ARR prompt (as in Figure 2) is paraphrased into five different versions [^4] Please refer to Appendix B for details on variants. by GPT-4o Hurst et al. (2024). Figure 3 shows the average performance of each method over 10 QA datasets, and the full results are presented in Table 12. As illustrated, the proposed ARR method, regardless of its prompt implementation, remains a consistent advantage over the baselines, substantiating that ARR is a general framework for question answering.